OpenStudy (anonymous):

What is theta of 6+6i if it is not 45degrees???? I'm kinda stuck. . .

OpenStudy (anonymous):

I know the polar form is r(costheta +sin i theta) but I really honestly dont know!!! I know its stupid but I've actually been at these questions for hours straight and I think my brain is fried! Whats even worse is that I have another one thats saying find r and theta of -8-8i as well and I'm just seeing 45degrees everywhere and its not being accepted by the damn computer system ......

OpenStudy (anonymous):

I think I'm gonna cry :(

OpenStudy (anonymous):

And theres no one around to hug me. :( There should be a warning before you start maths assignments telling you to make sure someone is near you at all times. . . .

OpenStudy (anonymous):

@Loser66 - WHERE DID YOU GO??????? D:

OpenStudy (anonymous):

it is 45degrees

OpenStudy (anonymous):

well that helps @dpasingh . . . .

OpenStudy (loser66):

@dpasingh please help her/ him know how to get it. It's better than give her/him answer. I let you finish your starting.

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

given z= 6+6i compare it with z=a+bi we have a= 6 and b= 6 Now \[\theta = \tan^{-1} (\frac{b}{a}) =\tan^{-1} (\frac{6}{6}) =\tan^{-1} (1) =\tan^{-1} (\tan 45^0) =45^0\] Thus θ= 45 degrees

OpenStudy (anonymous):

But i said that the answer was not 45 degrees !!! :/

OpenStudy (anonymous):

Your Ans is wrong, it is only 45 degrees.

OpenStudy (anonymous):

@HazelL

OpenStudy (anonymous):

@HazelL perhaps you was typing something.

OpenStudy (anonymous):

I love you @Loser66